OpenStudy (anonymous):

Designer Dolls, Inc. found that the number N of dolls sold varies directly with their advertising budget A and inversely with the price P of each doll. The company sold 5200 dolls when $26,000 was spent on advertising and the price of a doll was set at $30. Determine the number of dolls sold when the amount spent on advertising is increased to $52,000. Round to the nearest whole number. A. 5,200 dolls B. 1,723 dolls C. 3,447 dolls D. 10,400 dolls

OpenStudy (anonymous):

Are you going to Ashworth?

OpenStudy (anonymous):

yes, my last class

OpenStudy (anonymous):

part 1 or 2

OpenStudy (ganpat):

N = Number of dolls.. A = Advertisement , P = Price.. N directly A and N inversely to P \[N \alpha (A/P) ... = N = K(A/P)\] where K = constant of proportionality.. case 1: N = 5200, A = 26,000$ and P = 30$ N = K (A/P) 5200 = K (26000/30) K = 6.. Now, part 2, calculate N , when K = 6, p = 30 & A = 52000 so N = K (A/P) N = 6 (52000/30) N = 10400... option D...

OpenStudy (anonymous):

@Ganpat are you sure it's D????

OpenStudy (ganpat):

yup... calculation says so.. :) whats ur opinion on my approach ?? :D

OpenStudy (anonymous):

idk i'm horrible at math D: But we'll see when I turn my exam in!

OpenStudy (anonymous):

The answer is D

OpenStudy (ganpat):

thats bad.. From next time, we will practice maths before exams.. All the Best.. and m 100% sure :)

OpenStudy (anonymous):

thank you!

OpenStudy (ganpat):

W/c :)

OpenStudy (anonymous):

you were right btw(:

OpenStudy (ganpat):

so was i 100% sure.. so how much u scored ?